# Agnostic Sharpness-Aware Minimization

### v2

@@ 3.3 Our Solution
     Sample mini-batch $B\sim S$.
     Randomly split $B$ to $B^{t}$ and $B^{v}$.
     Compute $\tilde{\theta}_{l}^{v}=\theta_{l}+\rho_{2}\frac{\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)\|_{2}}$.
     Compute $\tilde{\theta}_{l}^{t}=\theta_{l}+\rho_{1}\frac{\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\|_{2}}-\rho_{2}\frac{\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)\|_{2}}$.
     Compute $\theta_{l+1}=\theta_{l}-\eta\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$.
@@ 4 Experiments
In this section, we present the results of various experiments to evaluate the effectiveness of our Agnostic SAM, including training from scratch, transfer learning on various datasets and learning with noisy labels. In all experiments, we compare our proposed method with SAM and SGD as baselines.
@@ 4.1 Image Classification From Scratch
We first conduct experiments on the CIFAR dataset using three architectures: WideResnet28x10, Pyramid101, and Densenet121. We compared the performance of our method with baseline models trained with the SGD and SAM optimizers. When replicating the experiments with SAM, we follow the recommended hyperparameters in the original paper, setting $\rho=0.05$ for CIFAR-10 and $\rho=0.1$ for CIFAR-100 experiments. For our AgnosticSAM approach, we consistently applied $\rho_{1}=2\rho_{2}$ across all experiments. Specifically, we set $\rho_{1}=0.1$, $\rho_{2}=0.05$ for experiments with CIFAR-10 and $\rho_{1}=0.2$, $\rho_{2}=0.1$ for experiments with CIFAR-100.
All experiments with SAM and AgnosticSAM are trained with 200 epochs with basic data augmentations (horizontal flip, padding by four pixels, and random crop) and an initial learning rate of 0.1 with a cosine learning schedule. Each experiment was conducted three times with different random seeds, and we reported the mean and standard deviation of the results in Table 1. The SGD results are referenced from [16].
Our proposed method consistently outperforms the baselines across various settings. Performance on the CIFAR-10 dataset is near the situation point, making it harder to get a notable improvement. However, our AgnosticSAM still manages to achieve slight enhancements in all cases. In experiments with the CIFAR-100 dataset, models are easier to get overfitting compared to CIFAR-10. In this case, our AgnosticSAM method consistently outperforms SAM, showing substantial improvements ranging from 1.17% to 1.67% in all experiments.
@@
SAM 96.72 $\pm$ 0.007 96.20 $\pm$ 0.134 91.16 $\pm$ 0.240
Agnostic-SAM 97.17 $\pm$ 0.014 96.30 $\pm$ 0.070 91.17 $\pm$ 0.148
CIFAR-100 SGD 81.20 $\pm$ 0.200 80.30 $\pm$ 0.300 -
SAM 82.69 $\pm$ 0.035 81.26 $\pm$ 0.636 68.09 $\pm$ 0.403
Agnostic-SAM 83.86 $\pm$ 0.077 82.93 $\pm$ 0.247 69.65 $\pm$ 0.226
Beyond the CIFAR dataset, we evaluate AgnosticSAM on multiple small-size datasets:
  • Oxford 102 Flower (Flower102) consists of 102 flower categories. Each class consists of between 40 and 258 images. We concatenate both the training set and validation set to use for the training process with 2,040 images in total and use the remaining 6,149 images for the testing set.
  • Stanford Cars contains 96 classes of cars with 16,185 images in total. The dataset is split into 8,144 training images and 8,041 testing images, where each class has been divided roughly in a 50-50 split.
  • FGVC-Aircraft consists of 10,200 images with 100 images for each of 102 different classes. This data is divided into 6,667 images for training and 3,533 images for testing.
  • Oxford IIIT Pets has 37 pet categories with roughly 200 images for each class. This dataset has 7,390 images in total and is split into 3,680 images for training and 3,710 images for testing.
Our experiments are conducted on three architectures: EfficientNet-B4, EfficientNet-B3, and ResNet18, all trained from scratch with basic data augmentation techniques (random resizing, cropping to a size of 224x224 pixels, horizontal flipping, changing brightness and contrast, and normalization). We ensure that the baselines are produced using the same hyperparameters as AgnosticSAM, including the learning rate, the number of training epochs, and the learning schedule. For experiments with SAM, we examine multiple values of $\rho\in\{0.02,0.05,0.1\}$ and report the highest accuracy achieved. In the case of AgnosticSAM, we examine with a fixed $\rho_{1}=0.1,\rho_{2}=0.05$, which is the same setting as experiments on the CIFAR-10 dataset. The results of these experiments are presented in Table 2.
It is notable that, in most cases, SAM achieves better accuracy than SGD after tuning with various values of $\rho$. However, performances seem to be unstable on the EfficientNet architectures. In contrast, AgnosticSAM consistently outperforms the baselines across all experiments, showing significant improvement gaps. This demonstrates the robustness and effectiveness of AgnosticSAM in enhancing model generalization under data limitations.
Table 2: Classification accuracy small-size datasets. All models are trained from scratch with basic data augmentation methods.
Method Flower102 Stanford Cars FGVC-Aircraft Oxford IIIT Pets
@@
SGD 69.56 82.27 79.57 46.33
SAM 68.58 81.74 81.22 61.16
Agnostic-SAM 76.86 85.15 84.70 66.48
EfficientNet-B3
SGD 71.33 81.26 80.92 54.87
SAM 73.93 81.31 81.79 65.74
Agnostic-SAM 75.39 83.53 85.90 72.91
ResNet-18
SGD 72.13 85.47 80.41 57.85
SAM 72.45 86.03 81.49 58.19
Agnostic-SAM 77.36 87.89 85.30 69.28
4.2 Transfer learning
Transfer learning is one of the most common techniques that enhances the learning process when dealing with small datasets. This method employs pre-trained models on a large dataset as the initial weights for fine-tuning on a smaller dataset to improve performance and accelerate the learning process. In this section, we further evaluate Agnostic-SAM in the transfer learning context using the ImageNet pre-trained models on EfficientNet-B4 to fine-tune three small-size datasets: Oxford Flowers-102, Stanford Cars, and FGVC-Aircraft. For each experiment, we fine-tune with a learning rate of 0.01 in 100 epochs and use $\rho=0.1$ for all experiments of SAM (as the accuracies tend to decrease when decreasing $\rho$), $\rho_{1}=2\rho_{2}=0.1$ for all experiments of AgnosticSAM. In Table 3, SAM performs more stable compared to the training from scratch setting discussed in Section 4.1, Agnostic-SAM achieves a noticeable improvement compared to baselines demonstrating its robustness across various experiment settings.
Table 3: Transfer learning on EfficientNet-B4 to small-size dataset with pre-trained weight on ImageNet dataset.
Dataset SGD SAM AgnosticSAM
Flower102 93.15 93.32 94.19
Stanford Cars 84.36 84.94 88.66
FGVC-Aircraft 80.86 81.79 85.09
@@ 4.3 Train With Noisy Label
In addition to mitigating data shifts between training and testing datasets, we evaluate the robustness of AgnosticSAM against noisy labels. While other methods may utilize various techniques to train with noisy labels, such as fine-tuning from pre-trained models or employing architectures specifically designed for noisy-label scenarios, we applied the standard training procedure to enhance the robustness of AgnosticSAM. Specifically, we adopt a classical noisy-label setting for CIFAR-10 and CIFAR-100, where a portion of the training set’s labels are randomly flipped with a fraction {0.1, 0.2, 0.3, 0.4}, and the testing set’s labels remain unchanged.
We conduct all experiments using the ResNet32 architecture, training from scratch for 200 epochs. The hyperparameters are consistent with those used in the CIFAR dataset experiments described in Section 4.1. Particularly, we set $\rho=0.05$ for CIFAR-10 and $\rho=0.1$ for CIFAR-100 with SAM experiments. For AgnosticSAM, we set $\rho_{1}=2\rho_{2}=0.1$ for CIFAR-10 and $\rho_{1}=2\rho_{2}=0.2$ for CIFAR-100. Each experiment is conducted three times using different random seeds, and we report the average and standard deviation of the results.
As shown in Table 4, our AgnosticSAM outperforms the other baselines by significant margins. As the noise fraction increases, the performance of the baselines downgrades considerably. Specifically, there is a 24.84% decrease in accuracy for the SGD optimizer and a 7.14% decrease for the SAM optimizer on the CIFAR-10 dataset when the noise fraction changes from 0.1 to 0.4. This figure is only 4.62% for our AgnosticSAM. Moreover, the performance gap between AgnosticSAM and the other baselines becomes significant when the noise fraction increases. This trend is consistent on the CIFAR-100 dataset, further demonstrating the stability and robustness of our method against label noise.
Table 4: Test accuracy on the clean test set for models trained on CIFAR-10 with noisy labels. Each experiment is conducted three times using different random seeds, and we report the average and standard deviation of the results.
@@
0.1 0.2 0.3 0.4
@@
SGD 89.47 $\pm$ 0.162 84.05 $\pm$ 1.103 73.66 $\pm$ 0.629 64.63 $\pm$ 0.537
SAM 93.04 $\pm$ 0.134 92.38 $\pm$ 0.148 90.14 $\pm$ 0.282 85.90 $\pm$ 0.084
Agnostic-SAM 93.71 $\pm$ 0.120 93.38 $\pm$ 0.233 91.98 $\pm$ 0.388 89.09 $\pm$ 0.537
Dataset CIFAR-100
SGD 67.47 $\pm$ 0.113 64.88 $\pm$ 0.657 58.38 $\pm$ 0.106 49.00 $\pm$ 1.527
SAM 70.76 $\pm$ 0.233 69.09 $\pm$ 0.212 64.72 $\pm$ 0.106 54.65 $\pm$ 0.516
Agnostic-SAM 71.61 $\pm$ 0.267 70.04 $\pm$ 0.155 66.72 $\pm$ 0.014 56.67 $\pm$ 0.261

### v3

@@ 3.3 Our Solution
     Sample mini-batch $B^{t}\sim S^{t}$, $B^{v}\sim S^{v}$.
     Compute $\tilde{\theta}_{l}^{v}=\theta_{l}+\rho_{2}\frac{\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\theta_{l}\right)\|_{2}}$
     $g_{v}\leftarrow\beta g_{v}+(1-\beta)\nabla_{\theta}\mathcal{L}_{B^{v}}\left(\tilde{\theta}_{l}^{v}\right)$
     Compute $\tilde{\theta}_{l}^{t}\leftarrow\theta_{l}+\rho_{1}\frac{\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)}{\|\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\theta_{l}\right)\|_{2}}-\rho_{2}\frac{g_{v}}{\|g_{v}\|_{2}}$.
     Compute $\theta_{l+1}\leftarrow\theta_{l}-\eta\nabla_{\theta}\mathcal{L}_{B^{t}}\left(\tilde{\theta}_{l}^{t}\right)$.
@@ 4 Experiments
In this section, we present the results of various experiments to evaluate the effectiveness of our Agnostic-SAM, including training from scratch, transfer learning on different dataset sizes, learning with noisy labels, and MAML setting. For all experiments of Agnostic-SAM, we consistently use a fixed value of momentum factor $\beta=0.9$ and mini-batch size of validation set $4|B^{v}|=|B^{t}|$. The effectiveness of these hyper-parameters on performance and training complexity will be discussed in Section 5.
@@ 4.1 Image Classification From Scratch
We first conduct experiments on ImageNet, Food101, and CIFAR datasets with standard image classification settings trained from scratch. The performance is compared with baseline models trained with the SGD, SAM, ASAM, and the integration of ASAM and Agnostic-SAM.
ImageNet dataset
We use ResNet18 and ResNet34 models for experiments on the ImageNet dataset, with an input size of $224\times 224$. For all experiments of Agnostic-SAM and its variations, we consistently set $\rho_{1}=2\rho_{2}=2\rho$, where $\rho$ represents the perturbation radius for the respective SAM method. Specifically, in this experiment, we set $\rho=0.1$, $\rho_{1}=0.2$, and $\rho_{2}=0.1$. The models are trained for 200 epochs with basic data augmentations (random cropping, horizontal flipping, and normalization). We use an initial learning rate of 0.1, a batch size of 2048 for the training set, and 512 for the validation set, following a cosine learning schedule across all experiments in this paper. We extend this experiment to the mid-sized Food101 dataset using the same settings, except for a batch size of $128$ for the training set and $32$ for the validation set. Performance results are detailed in Table 1.
Table 1: Classification accuracy on the ImageNet and Food101 datasets. All models are trained from scratch with 200 epochs.
Dataset Method Resnet18 Resnet34
Top-1 Top-5 Top-1 Top-5
ImageNet SAM 62.46 84.19 63.73 84.95
Agnostic-SAM 63.64 85.22 65.89 86.84
Food101 SAM 73.15 89.85 73.87 90.84
Agnostic-SAM 73.45 90.35 74.47 91.27
CIFAR dataset
We used three architectures: WideResnet28x10, Pyramid101, and Densenet121 with an input size of $32\times 32$ for CIFAR datasets. To replicate the baseline experiments, we followed the hyperparameters provided in the original papers. Specifically, for CIFAR-100, we set $\rho=0.1$, $\rho_{1}=0.2$, and $\rho_{2}=0.1$, and for CIFAR-10, we used $\rho=0.05$, $\rho_{1}=0.1$, and $\rho_{2}=0.05$. The same procedure and settings were applied to ASAM and Agnostic-ASAM, with the perturbation radius $\rho$ for ASAM being 10 times larger than that of the SAM method. Other training configurations are consistent with those used in the ImageNet experiments, except for data augmentations (horizontal flipping, four-pixel padding, and random cropping). Results are reported in Tables 4.1, while the SGD results are referenced from Foret et al. (2021).
Our proposed method consistently outperforms the baselines across various settings. On both ImageNet and Food101, it significantly surpasses the baselines, with a notable improvement in both Top-1 and Top-5 accuracy. For CIFAR-10, performance is close to the saturation point, making further improvements challenging. Nevertheless, Agnostic-SAM achieves slight enhancements across all cases. On CIFAR-100, where models are more prone to overfitting compared to CIFAR-10, Agnostic-SAM still delivers competitive results.
@@
SAM 96.87 $\pm$ 0.027 96.17 $\pm$ 0.174 91.28 $\pm$ 0.241
Agnostic-SAM 96.88 $\pm$ 0.007 96.47 $\pm$ 0.219 91.31 $\pm$ 0.707
[…]
Agnostic-ASAM 97.15 $\pm$ 0.063 96.73 $\pm$ 0.261 92.02 $\pm$ 0.000
4.2 Transfer learning
In this subsection, we further evaluate Agnostic-SAM in the transfer learning setting using the ImageNet pre-trained models to fine-tune both small-size, mid-size, and large-size datasets. All initialized weights are available on the Pytorch library.
Table 3: Transfer learning on ImageNet with Resnet models.
Model Top-1 Acc Top-5 Acc
SAM Agnostic-SAM SAM Agnostic-SAM
Resnet18 70.52 70.88 89.60 89.94
Resnet34 73.06 73.84 91.29 91.81
Resnet50 75.17 75.91 92.58 92.83
First, we conduct experiments on ImageNet by using three models from the Resnet family. These base models are both pre-trained on ImageNet by SGD and then fine-tuned for 50 epochs by SAM or Agnostic-SAM with a learning rate of $0.01$. We $\rho=0.05$ for SAM, and $\rho_{1}=2\rho_{2}=0.1$ for Agnostic-SAM and basic augmentation techniques, which are the same as training from the scratch setting. Results reported in Table 3 show that our methods outperform baselines with a significant gap in both top-1 and top-5 accuracies.
Next, we examine this setting on small and mid-sized datasets on three models of the EfficientNet family. We fine-tune with a learning rate of 0.05 in 50 epochs and use $\rho=0.1$ for all experiments of SAM (as accuracies tend to decrease when reducing $\rho$), $\rho_{1}=2\rho_{2}=0.1$ for all experiments of Agnostic-SAM. In Table 4, Agnostic-SAM achieves a noticeable improvement compared to most of the baselines on all small-size, mid-size, and large-size datasets, demonstrating its robustness and stability across various experiment settings.
Table 4: Transfer learning accuracy of small and medium datasets. All models are fine-tuned from pre-trained weights on ImageNet.
Dataset Top-1 Acc Top-5 Acc
SGD SAM Agnostic-SAM SGD SAM Agnostic-SAM
EfficientNet-B2
Stanford Cars 89.14 $\pm$ 0.11 89.68 $\pm$ 0.17 90.34 $\pm$ 0.07 97.60 $\pm$ 0.20 98.04 $\pm$ 0.07 98.24 $\pm$ 0.09
FGVC-Aircraft 85.83 $\pm$ 0.23 86.25 $\pm$ 0.36 87.27 $\pm$ 0.27 95.72 $\pm$ 0.02 95.87 $\pm$ 0.06 96.05 $\pm$ 0.03
Oxford IIIT Pets 92.17 $\pm$ 0.19 92.34 $\pm$ 0.11 92.58 $\pm$ 0.17 99.23 $\pm$ 0.02 99.35 $\pm$ 0.02 99.35 $\pm$ 0.07
Flower102 95.06 $\pm$ 0.01 95.22 $\pm$ 0.14 95.56 $\pm$ 0.10 99.08 $\pm$ 0.18 99.11 $\pm$ 0.19 99.27 $\pm$ 0.02
Food101 83.50 $\pm$ 0.01 85.12 $\pm$ 0.07 85.51 $\pm$ 0.02 96.10 $\pm$ 0.32 96.83 $\pm$ 0.08 97.14 $\pm$ 0.00
Country211 11.94 $\pm$ 0.14 12.48 $\pm$ 0.03 13.28 $\pm$ 0.00 23.70 $\pm$ 0.13 25.49 $\pm$ 0.07 26.95 $\pm$ 0.16
EfficientNet-B3
Stanford Cars 89.01 $\pm$ 0.19 89.40 $\pm$ 0.09 90.09 $\pm$ 0.14 97.73 $\pm$ 0.21 98.03 $\pm$ 0.07 98.13 $\pm$ 0.01
FGVC-Aircraft 84.88 $\pm$ 0.08 85.19 $\pm$ 0.11 85.99 $\pm$ 0.25 95.53 $\pm$ 0.12 95.67 $\pm$ 0.00 96.08 $\pm$ 0.10
Oxford IIIT Pets 92.68 $\pm$ 0.25 92.58 $\pm$ 0.02 92.75 $\pm$ 0.19 99.00 $\pm$ 0.01 99.19 $\pm$ 0.05 99.20 $\pm$ 0.11
Flower102 94.59 $\pm$ 0.10 94.73 $\pm$ 0.14 95.16 $\pm$ 0.26 98.95 $\pm$ 0.08 99.12 $\pm$ 0.16 99.18 $\pm$ 0.07
Food101 83.75 $\pm$ 0.12 85.79 $\pm$ 0.13 86.17 $\pm$ 0.13 96.22 $\pm$ 0.02 97.12 $\pm$ 0.00 97.38 $\pm$ 0.00
Country211 12.96 $\pm$ 0.01 13.38 $\pm$ 0.09 13.63 $\pm$ 0.05 26.11 $\pm$ 0.56 25.78 $\pm$ 0.08 26.71 $\pm$ 0.26
@@ 4.2 Transfer learning
Stanford Cars 84.72 $\pm$ 0.04 85.08 $\pm$ 0.16 85.79 $\pm$ 0.32 96.41 $\pm$ 0.07 96.45 $\pm$ 0.01 96.77 $\pm$ 0.00
FGVC-Aircraft 79.95 $\pm$ 0.61 79.96 $\pm$ 0.04 80.80 $\pm$ 0.51 94.87 $\pm$ 0.08 94.65 $\pm$ 0.08 94.95 $\pm$ 0.01
Oxford IIIT Pets 91.89 $\pm$ 0.13 92.02 $\pm$ 0.23 92.02 $\pm$ 0.00 99.28 $\pm$ 0.10 99.43 $\pm$ 0.07 99.44 $\pm$ 0.02
Flower102 92.73 $\pm$ 0.04 93.02 $\pm$ 0.14 93.03 $\pm$ 0.16 98.49 $\pm$ 0.07 98.68 $\pm$ 0.02 98.59 $\pm$ 0.05
Food101 84.55 $\pm$ 0.14 86.13 $\pm$ 0.06 86.15 $\pm$ 0.44 96.31 $\pm$ 0.03 97.07 $\pm$ 0.01 97.22 $\pm$ 0.02
Country211 14.63 $\pm$ 0.09 14.80 $\pm$ 0.13 15.10 $\pm$ 0.16 27.60 $\pm$ 0.00 29.09 $\pm$ 1.77 28.38 $\pm$ 0.14
@@ 4.3 Train With Noisy Label
In addition to mitigating data shifts between training and testing datasets, we evaluate the robustness of Agnostic-SAM against noisy labels on standard training procedure. Specifically, we adopt a classical noisy-label setting for CIFAR-10 and CIFAR-100, in which a portion of the training set’s labels are symmetrically flipped with noise fractions {0.2, 0.4, 0.6, 0.8}, while the testing set’s labels remain unchanged.
All experiments are conducted using the ResNet32 architecture, with models trained from scratch for 200 epochs. The batch size is set to 512 for the training set and 128 for the validation set. Following Li et al. (2024) and Foret et al. (2021), we set $\rho=0.1$ and $\rho_{1}=2\rho_{2}=0.2$ for SAM, FSAM, and Agnostic-SAM, $\rho_{1}=2\rho_{2}=2\rho=2.0$ for ASAM and Agnostic-ASAM when training with all noise levels, except for the 80%, where we reduce the perturbation radius by half to ensure more stable convergence. In line with Li et al. (2024), we apply additional cutout techniques along with the basic augmentations outlined in Section 4.1. Each experiment is repeated three times with different random seeds, and we report the average and standard deviation of the results in Table 4.3. Note that training with SGD is prone to overfitting as the number of epochs increases. Therefore, we present the best results for SGD training at both 200 and 400 epochs.
Table 5: Results under label noise on CIFAR dataset with ResNet32. Each experiment is conducted three times using different random seeds, and we report the average and standard deviation of the results.
@@
0.2 0.4 0.6 0.8
Dataset CIFAR-100
SGD 66.22 $\pm$ 0.355 59.26 $\pm$ 0.045 46.77 $\pm$ 0.020 26.49 $\pm$ 0.640
SAM 66.16 $\pm$ 0.721 59.95 $\pm$ 0.622 50.81 $\pm$ 0.353 24.26 $\pm$ 1.209
FSAM 65.73 $\pm$ 0.219 58.96 $\pm$ 0.381 49.36 $\pm$ 1.103 25.92 $\pm$ 1.173
Agnostic-SAM 66.64 $\pm$ 0.657 61.13 $\pm$ 0.636 52.26 $\pm$ 0.502 27.66 $\pm$ 1.265
[…]
Agnostic-ASAM 67.38 $\pm$ 0.106 62.72 $\pm$ 0.304 54.58 $\pm$ 0.572 32.77 $\pm$ 0.388
@@
SGD 89.98 $\pm$ 0.070 84.83 $\pm$ 0.085 75.06 $\pm$ 0.385 54.47 $\pm$ 1.265
SAM 91.26 $\pm$ 0.007 88.19 $\pm$ 1.060 83.43 $\pm$ 0.622 61.69 $\pm$ 0.289
FSAM 91.35 $\pm$ 0.318 87.58 $\pm$ 0.353 82.78 $\pm$ 2.057 58.09 $\pm$ 2.276
Agnostic-SAM 92.38 $\pm$ 0.007 90.20 $\pm$ 0.318 85.33 $\pm$ 0.268 70.02 $\pm$ 0.403
[…]
Agnostic-ASAM 92.06 $\pm$ 0.367 90.01 $\pm$ 0.282 86.09 $\pm$ 0.657 73.25 $\pm$ 0.353
4.4 Experiments on meta-learning
The concept of Agnostic-SAM is inspired by the agnostic approach in the MAML setting, where the meta-model is optimized on the meta-training set but aims to minimize loss on the validation set. The key difference is that Agnostic-SAM uses the gradient from the validation set as an indicator to close the generalization gap between the training and testing sets. Despite this difference, both approaches share the same underlying principle, making it reasonable to expect that applying Agnostic-SAM in the MAML setting will result in improved generalization performance.
Table 6: Meta-learning results on Mini-Imagenet dataset. All baseline results are taken from Abbas et al. (2022)
Method Accuracy
5 ways 1 shot 5 ways 5 shots
MAML 47.13 62.20
SHARP-MAML 49.72 63.18
Agnostic-SAM 50.08 64.29
We compare our approach to standard MAML and Sharp-MAML (Abbas et al., 2022), which also addresses the loss-landscape flatness in bilevel models. The experiments follow the setup from Abbas et al. (2022), specifically using the Sharp-MAMLlow variation, which focuses on minimizing the sharpness of meta-models trained on the meta-training set. Note that during the testing phase of MAML, only the meta-training set is used for a few update steps of the meta-model; and our Agnostic-SAM approach incorporates both the training and validation sets in the meta-model training process. Ideally, both the meta-training and meta-validation sets should be utilized to minimize the lower-level loss during training. However, this could introduce inconsistencies between the training and testing phases, potentially degrading performance during testing. To avoid this issue, we duplicate the meta-training set and use it as a validation set to minimize the lower-level loss of the meta-model, applying this procedure consistently during both the training and testing phases.
As with other experiments, we set $\rho_{1}=2\rho_{2}=2\rho$, with $\rho$ as the perturbation radius for Sharp-MAMLlow, and report the results in Tables 6 and 7. Our method consistently outperforms most baselines with significant improvements, demonstrating the effectiveness of Agnostic-SAM and its flexibility across various settings.
Table 7: Meta-learning results on Omniglot dataset. All baseline results are taken from Abbas et al. (2022)
Method Accuracy
20 ways 1 shot 20 ways 5 shots
MAML 91.77 96.16
SHARP-MAML 92.89 96.59
Agnostic-SAM 92.66 97.28